# Sharp one-point estimates and Minkowski content for the scaling limit of three-dimensional loop-erased random walk

### v1

@@ 4 Sharp one-point function estimates
In this subsection, we deal with the asymptotics of the probability that a LERW passes a ball and prove Theorem 1.4 on the one-point function for the scaling limit. The idea of proof is somewhat similar to how we derive Theorem 1.3 and [19, Theorem 1.1], however some special care is needed in decomposing the path, and a new decoupling technique is involved.
@@
We now comment on the proof of Theorem 1.4. The strategy consists of checking that the one-point function in Theorem 1.4 is a multiple of the (discrete) one-point function in Theorem 1.3. We first compare hitting probabilities for balls around $x\in\mathbb{D}$ and the reference point $\hat{x}$.
@@
$\frac{P(\eta_{m}\cap B(\hat{x},s)\neq\emptyset)}{P((\hat{x})_{m}\in\eta_{m})}=\frac{P(\eta_{m}\cap B(x,s)\neq\emptyset)}{P(x_{m}\in\eta_{m})}\Big{[}1+O\big{(}d_{x}^{-c}m^{-\delta}\big{)}\Big{]}$ (4.15)
@@
We omit the proof of Proposition 4.7 as it follows from the same coupling technique as the proof of Proposition 4.5 above, with the only extra ingredient being the rotation invariance of the scaling limit of 3D LERW.
@@
We now work towards the proof of Proposition 4.5. We begin introducing notation that will be used for the rest of this section.
@@
    $\partial_{i}B(x_{0},s)=\Bigg{\{}y\in m^{-1}\mathbb{Z}^{3}\cap B(x_{0},s)\ \Bigg{|}\ \begin{array}[]{c}\text{There exists }\\ z\in m^{-1}\mathbb{Z}^{3}\cap B(x_{0},s)^{c}\\ \text{ such that }|y-z|=m^{-1}\end{array}\Bigg{\}}$ (4.16)
    be the (discrete) inner boundary of $B(x_{0},s)$.
@@
    be the first time that $S^{1}$ hits the origin.
@@
  • Let
    $u_{1}=\max\Big{\{}j\in[0,u_{2}]\ \Big{|}\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}(j)\in B(x_{0},s)\Big{\}}$ (4.20)
    be the last time that the loop-erasure of $S^{1}$ stays in $B(x_{0},s)$.
@@
By definition of the loop-erasure, the LHS of (4.13) is equal to the following probability:
$\sum_{y\in\partial_{i}B(x_{0},s)}P^{y,y}_{1,2}\Bigg{(}\begin{array}[]{c}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\Bigg{)}.$ (4.21)
Here we used notation introduced above, between (4.16) and (4.20).
We would want to say that the probability in the sum of (4.21) actually does not depend on $y\in\partial_{i}B(x_{0},s)$. However, unfortunately, it is not the case. To deal with this issue, we use a coupling method as follows.
We first try to change the definition of $u_{1}$, in (4.20), by replacing the definition of $u_{2}$, in (4.19), with some earlier time in the definition. Let $\alpha=2-\beta\in[1/3,1)$. We set
@@
$v_{1}=\inf\Big{\{}j\geq 0\ \Big{|}\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}(j)\notin B\big{(}x_{0},s^{\delta_{0}}\big{)}\Big{\}}\text{ and }$
$u_{1}^{\prime}=\max\Big{\{}j\in[0,v_{1}]\ \Big{|}\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}(j)\in B(x_{0},s)\Big{\}}.$
Note that
$P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ u_{1}\neq u_{1}^{\prime},\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}\leq Cs^{-\delta_{0}}m^{-2}.$
On the other hand, it holds that
$P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)\asymp s^{-1+\alpha}m^{-2}.$
Therefore, we have for each term in (4.21) that
$P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
$=P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ u_{1}=u_{1}^{\prime},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)$
$\ \ \ +P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ u_{1}\neq u_{1}^{\prime},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)$
$=P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ u_{1}=u_{1}^{\prime},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)$
$\ \ \ +P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)O\big{(}s^{\delta_{0}}\big{)}$
$=P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)$
$\ \ \ +P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)O\big{(}s^{\delta_{0}}\big{)},$
and equivalently
$P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
$=P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
×$\ \ \ \ \ \ \ \ \ \times\Big{[}1+O\big{(}s^{\delta_{0}}\big{)}\Big{]}.$
Thus, we can replace $u_{1}$ by $u_{1}^{\prime}$ which is easier to analyze. Indeed, if we let
$t_{2}=\inf\Big{\{}j\geq 0\ \Big{|}\ S^{2}(j)\notin B\big{(}x_{0},s^{\delta_{0}}\big{)}\Big{\}}$ (4.23)
be the first time that $S^{2}$ leaves the ball $B\big{(}x_{0},s^{\delta_{0}}\big{)}$, then by the coupling argument it follows that for all $y,z\in\partial_{i}B(x_{0},s)$
$\frac{P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{z,z}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}}$
$=\frac{P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},v_{1}]\cap S^{2}[1,t_{2}]=\emptyset,\\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)}{P^{z,z}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},v_{1}]\cap S^{2}[1,t_{2}]=\emptyset,\\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)}\Big{[}1+O\big{(}s^{\delta_{1}}\big{)}\Big{]},$
for some universal constant $\delta_{1}>0$. However, it is easy to see that for all $y,z\in\partial_{i}B(x_{0},s)$
@@
Thus, if we let
$b=\inf\Big{\{}j\geq 0\ \Big{|}\ \text{LE}\big{(}S^{1}[0,\infty)\big{)}(j)\notin B\big{(}x_{0},s^{\delta_{0}}\big{)}\Big{\}}\text{ and }$
$a=\max\Big{\{}j\in[0,b]\ \Big{|}\ \text{LE}\big{(}S^{1}[0,\infty)\big{)}(j)\in B(x_{0},s)\Big{\}},$ (4.24)
then it follows from Proposition 4.4 of [21] that for all $y,z\in\partial_{i}B(x_{0},s)$
$\frac{P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{z,z}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1}^{\prime},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}}$
$=\frac{P^{y,y}_{1,2}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{z,z}_{1,2}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}\Big{[}1+O\big{(}s^{\delta_{1}}\big{)}\Big{]}.$ (4.25)
We write $\gamma^{1}=\text{LE}\big{(}S^{1}[0,\infty)\big{)}$ for the infinite LERW in $m^{-1}\mathbb{Z}^{3}$. For $L>1$, we let
$t_{2,L}=\inf\Big{\{}j\geq 0\ \Big{|}\ S^{2}(j)\notin B\Big{(}x_{0},s+s^{L}\Big{)}\Big{\}}.$
It is easy to see that
$P^{y}\left(\begin{array}[]{c}\text{diam}\big{(}S^{2}[0,t_{2,L}]\big{)}\geq s^{L/2},\\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)\leq Cm^{-1}s^{-L}\exp\big{\{}-cs^{-L/2}\big{\}}$
$=P^{y,y}_{1,2}\left(\begin{array}[]{c}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\end{array}\right)O\Big{(}\exp\big{\{}-cs^{-L/4}\big{\}}\Big{)}.$
@@
$P^{y,y}_{1,2}$ $\left(\begin{array}[]{c}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset,\ \text{diam}\big{(}S^{2}[0,t_{2,L}]\big{)}\leq s^{L/2}\end{array}\right)$
$=P^{y,y}_{1,2}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
×$\ \ \ \ \ \ \ \ \ \times\Big{[}1+O\Big{(}\exp\big{\{}-cs^{-L/4}\big{\}}\Big{)}\Big{]}.$
@@
$\Big{\{}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset\Big{\}}\text{ by }\Big{\{}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[t_{2,L},t_{2}]=\emptyset\Big{\}}.$
@@
$P^{y}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[0,b]\cap B\big{(}x_{0},s-s^{L/4}\big{)}=\emptyset\Big{)}\leq Cs^{\delta_{2}L},$
@@
$P^{y,y}_{1,2}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset,$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{LE}\big{(}S^{1}[0,\infty)\big{)}[0,b]\cap B\big{(}x_{0},s-s^{L/4}\big{)}\neq\emptyset\Big{)}$
$=P^{y,y}_{1,2}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}\Big{[}1+O\big{(}s\big{)}\Big{]}.$
@@
$\left\{\begin{array}[]{c}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2,L}]\neq\emptyset,\\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset,\ \text{diam}\big{(}S^{2}[0,t_{2,L}]\big{)}\leq s^{L/2}\end{array}\right\}$
and
$\Big{\{}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[0,b]\cap B\big{(}x_{0},s-s^{L/4}\big{)}\neq\emptyset\Big{\}}$

### v2

@@ 4 Sharp one-point function estimates
In this subsection, we deal with the asymptotics of the probability that a LERW passes through a ball and prove Theorem 1.4, which concerns the one-point function for the scaling limit. The general strategy of the proof is somewhat similar to that used in Theorem 1.3 and in [20, Theorem 1.1]. However, some special care is needed in the decomposition of the path, and a new decoupling technique plays a crucial role.
@@
We now comment on the proof of Theorem 1.4. The strategy consists of checking that the one-point function appearing in Theorem 1.4 is a constant multiple of the discrete one-point function established in Theorem 1.3. We first compare hitting probabilities for balls centered at a point $x\in\mathbb{D}$ and at the reference point $\hat{x}$.
@@
$\frac{P\Big(\eta_{m}\cap B(\hat{x},s)\neq\emptyset\Big)}{P\Big((\hat{x})_{m}\in\eta_{m}\Big)}=\frac{P\Big(\eta_{m}\cap B(x,s)\neq\emptyset\Big)}{P\Big(x_{m}\in\eta_{m}\Big)}\Big[1+O\big(d_{x}^{-c}m^{-\delta}\big)\Big]$ (4.18)
@@
We omit the proof of Proposition 4.7 as it follows from the same coupling technique as the proof of Proposition 4.5, with the only extra ingredient being the rotation invariance of the scaling limit of 3D LERW.
@@
We now work towards the proof of Proposition 4.5. First, recall that $S_{m}$ is a simple random walk on $m^{-1}\mathbb{Z}^{3}$ starting at 0 and stopped on its first exit of $\mathbb{D}_{m}$, and $\eta_{m}={\rm LE}(S_{m})$. Let us introduce additional notation that will be used for the rest of this section.
@@
    $\partial_{i}B(\hat{x},s)=\left\{y\in m^{-1}\mathbb{Z}^{3}\cap B(\hat{x},s)\ \Big|\ \begin{array}[]{c}\text{There exists }\\ z\in m^{-1}\mathbb{Z}^{3}\cap B(\hat{x},s)^{c}\\ \text{ such that }|y-z|=m^{-1}\end{array}\right\}$ (4.19)
    be the (discrete) inner boundary of $B(\hat{x},s)$.
@@
    be the first time that $S^{1}$ hits the origin and write
    $\gamma^{0}={\rm LE}(S^{1}[0,t^{1}_{0}])$
    for short.
@@
  • Let
    $\xi=\max\Big\{j\geq 0\ \Big|\ S(j)\in B(\hat{x},s)\Big\}$ (4.22)
    be the last time that the simple random walk $S_{m}$ stays in $B(\hat{x},s)$.
  • Let
    $u_{1}=\max\Big\{j\in[0,u_{2}]\ \Big|\ \gamma^{0}(j)\in B(\hat{x},s)\Big\}$ (4.23)
    be the last time that the loop-erasure of $S^{1}$ stays in $B(\hat{x},s)$.
@@
Let $y\in\partial_{i}B(\hat{x},s)$. If the random walk $S_{m}$ hits the vertex $y$, note that (up to time reversal) $S_{m}$ is equivalent to the concatenation of two independent simple random walks starting at $y$, one going forward in time, $S^{2}$, and the other going backward $S^{1}[0,t^{1}_{0}]$, conditioned on the event that $S^{1}$ hits 0 before exiting $\mathbb{D}$, i.e. $t^{1}_{0}<T^{1}$. If we further assume that the last visit to $B(\hat{x},s)$ occurs at time $\xi$ with $S_{m}(\xi)=y$, we obtain:
$P\left(S_{m}(\xi)=y\right)=P^{y,y}_{1,2}\left(t^{1}_{0}<T^{1},S^{2}[1,T^{2}]\cap B(\hat{x},s)=\emptyset\right).$
By definition of the loop-erasure, writing
$S^{2}_{+}=S^{2}[1,T^{2}]$
for short, we have
$P\left(S_{m}(\xi)=y,\eta_{m}\cap B(\hat{x},s)\neq\emptyset\right)=P^{y,y}_{1,2}\Bigg(\begin{array}[]{c}t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\end{array}\Bigg).$ (4.25)
Applying a last-exit decomposition and using (4.25), we obtain the following expression for the left-hand side of (4.16):
$P\left(\eta_{m}\cap B(\hat{x},s)\neq\emptyset\right)$ $=\sum_{y\in\partial_{i}B(\hat{x},s)}P\left(S_{m}(\xi)=y,\eta_{m}\cap B(\hat{x},s)\neq\emptyset\right)$
$=\sum_{y\in\partial_{i}B(\hat{x},s)}P^{y,y}_{1,2}\Bigg(\begin{array}[]{c}t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\end{array}\Bigg).$ (4.28)
Here, we use the notation introduced above, between (4.19) and (4.24).
We would like to say that the probability in the sum of (4.2) actually does not depend on $y\in\partial_{i}B(\hat{x},s)$. However, unfortunately, this is not the case. To deal with this issue, we use a coupling method as follows.
First, we will change the length of the loop-erasure of $S^{1}$ in the events appearing in (4.2). We will replace $u_{1}$ with the time $u_{1}^{\prime}$ as defined below in (4.2). Let $\alpha=2-\beta\in[1/3,1)$. We set
@@
$v_{1}=\inf\Big\{j\geq 0\ \Big|\ \gamma^{0}(j)\notin B\big(\hat{x},s^{\delta_{0}}\big)\Big\}\text{ and }$
$u_{1}^{\prime}=\max\Big\{j\in[0,v_{1}]\ \Big|\ \gamma^{0}(j)\in B(\hat{x},s)\Big\}.$ (4.30)
The difference between $u_{1}$ and $u_{1}^{\prime}$ is that $u_{1}$ considers the time interval $0,u_{2}]=[0,\text{len}(\gamma^{0})]$ (with $u_{2}$ is as defined in (4.24)), while $u_{1}^{\prime}$ considers a possibly shorter time interval $0,v_{1}]$. See Figure 2 for a graphical reference.
Figure 2: An illustration of the event $\{u_{1}\neq u_{1}^{\prime}\}$.
Note that the event $\{u_{1}\neq u_{1}^{\prime}\}$ is contained in the event that the LERW path $\gamma^{0}[v_{1},u_{2}]$intersects the ball $B(\hat{x},s)$. Then, by Lemma 4.4 and [13, Proposition 1.5.10]
$P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ u_{1}\neq u_{1}^{\prime},\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)\leq Cs^{-\delta_{0}}m^{-2},$ (4.31)
and by adding [28, Proposition 3.2.2] for the lower bound
$P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)\asymp s^{-1+\alpha}m^{-2},$ (4.32)
where we have used $\alpha<1$ to simplify the exponents. For each term in (4.2),
${\rm I}:=\;$ $P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)$
$=P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ u_{1}=u_{1}^{\prime},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\end{array}\right)$
$\ \ \ +P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ u_{1}\neq u_{1}^{\prime},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\end{array}\right)$
$=:{\rm II}+{\rm III}.$
By (4.31) and (4.32), one has that
${\rm III}=O\big(s^{\delta_{0}}\big)\cdot{\rm I},$
while it is easy to see that
${\rm II}\leq P^{y,y}_{1,2}\left(\begin{array}[]{c}t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\end{array}\right)\leq{\rm I},$
which implies
$P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)$
$=P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)\Big[1+O\big(s^{\delta_{0}}\big)\Big].$
Thus, we can replace $u_{1}$ by $u_{1}^{\prime}$, which is easier to analyze. Indeed, if we let
$t_{2}=\inf\Big\{j\geq 0\ \Big|\ S^{2}(j)\notin B\big(\hat{x},s^{\delta_{0}}\big)\Big\}$ (4.33)
be the first time that $S^{2}$ leaves the ball $B\big(\hat{x},s^{\delta_{0}}\big)$, and write
$S^{2}_{-}=S^{2}[1,t_{2}]$
for short. Then, by the coupling argument as stated in Lemma 3.3 of [22] (see also Lemma 5.4 of [22] for an analogous estimate concerning one-point function), it follows that for all $y,z\in\partial_{i}B(\hat{x},s)$
$\frac{P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)}{P^{z,z}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)}$
$\qquad=\frac{P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},v_{1}]\cap S^{2}_{-}=\emptyset,S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}{P^{z,z}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},v_{1}]\cap S^{2}_{-}=\emptyset,S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}\Big[1+O\big(s^{\delta_{0}}\big)\Big].$ (4.34)
However, it is easy to see that for all $y,z\in\partial_{i}B(\hat{x},s)$
@@
Recall that $\gamma^{0}$ denotes the loop-erasure of a simple random walk starting at $y\in B(\hat{x},s)$ and stopped upon hitting the origin 0. The segment $\gamma^{0}[u_{1}^{\prime},u_{2}]$ corresponds to the portion of this LERW between its last visit to the boundary of $B(\hat{x},s)$ (before exiting $B(\hat{x},s^{\delta_{0}})$) and the origin. In (4.34), we compared $\gamma^{0}[u_{1}^{\prime},u_{2}]$ with a shorter segment, $\gamma^{0}[u_{1}^{\prime},v_{1}]$, which is stopped upon leaving $B(\hat{x},s^{\delta^{0}})$ for the first time. We now proceed to compare $\gamma^{0}[u_{1}^{\prime},v_{1}]$ with a corresponding segment of an infinite LERW, which is more amenable to analysis. This comparison also extends to $\gamma^{0}[u_{1}^{\prime},u_{2}]$.
We write
$\gamma^{\infty}=\text{LE}\big(S^{1}[0,\infty)\big)$
for the infinite LERW in $m^{-1}\mathbb{Z}^{3}$. If we let
$b=\inf\Big\{j\geq 0\ \Big|\ \gamma^{\infty}(j)\notin B\big(\hat{x},s^{\delta_{0}}\big)\Big\}\text{ and }$
$a=\max\Big\{j\in[0,b]\ \Big|\ \gamma^{\infty}(j)\in B(\hat{x},s)\Big\},$ (4.35)
then it follows from (4.34) and [23, Proposition 4.4] that for all $y,z\in\partial_{i}B(\hat{x},s)$
$\frac{P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)}{P^{z,z}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1}^{\prime},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)}$
$\qquad=\frac{P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}{P^{z,z}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}\Big[1+O\big(s^{\delta_{1}}\big)\Big].$ (4.36)
For $L>1$, we let
$t_{2,L}=\inf\Big\{j\geq 0\ \Big|\ S^{2}(j)\notin B\Big(\hat{x},s+s^{L}\Big)\Big\}.$
By classical escape probability estimates for SRW, as in [13, Proposition 1.5.10], and the asymptotics of non-intersection probabilities of LERW and SRW from Theorem 4.3, one has that
$P^{y}\Big(\text{diam}\big(S^{2}[0,t_{2,L}]\big)\geq s^{L/2},S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)\leq Cm^{-1}s^{-L}O\big(s^{L/2}\big)$
$=P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)O\Big(s^{-L/4}\Big).$
@@
$P^{y,y}_{1,2}$ $\left(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,S^{2}_{-}\cap B(\hat{x},s)=\emptyset,\ \text{diam}\big(S^{2}[0,t_{2,L}]\big)\leq s^{L/2}\right)$
$=P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)\Big[1+O\Big(s^{-L/4}\Big)\Big].$
@@
$\Big\{\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset\Big\}\qquad\text{ by }\qquad\Big\{\gamma^{\infty}[a,b]\cap S^{2}[t_{2,L},t_{2}]=\emptyset\Big\}.$
@@
$P^{y}\Big(\gamma^{\infty}[0,b]\cap B\big(\hat{x},s-s^{L/4}\big)=\emptyset\Big)\leq Cs^{\delta_{2}L},$
@@
$P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset,\gamma^{\infty}[0,b]\cap B\big(\hat{x},s-s^{L/4}\big)\neq\emptyset\Big)$
$=P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)\Big[1+O\big(s\big)\Big].$
@@
$\left\{\begin{array}[]{c}\gamma^{\infty}[a,b]\cap S^{2}[1,t_{2,L}]\neq\emptyset,\\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset,\ \text{diam}\big(S^{2}[0,t_{2,L}]\big)\leq s^{L/2}\end{array}\right\}\mbox{ and }\Big\{\gamma^{\infty}[0,b]\cap B\big(\hat{x},s-s^{L/4}\big)\neq\emptyset\Big\}$